### 11 Mar 2026 v2

# Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper

Abstract
We present improved lower bounds for five classical Ramsey numbers: $\mathbf{R}(3,13)$ is increased from $60$ to $61$, $\mathbf{R}(3,18)$ from $99$ to $100$, $\mathbf{R}(4,13)$ from $138$ to $139$, $\mathbf{R}(4,14)$ from $147$ to $148$, and $\mathbf{R}(4,15)$ from $158$ to $159$. These results were achieved using AlphaEvolve, an LLM-based code mutation agent. Beyond these new results, we successfully recovered lower bounds for all Ramsey numbers known to be exact, and matched the best known lower bounds across many other cases. These include bounds for which previous work does not detail the algorithms used.
Virtually all known Ramsey lower bounds for “small” numbers are derived computationally, with bespoke search algorithms each delivering a handful of results. AlphaEvolve is a single meta-algorithm yielding search algorithms for all of our results.
1 Introduction
Ramsey numbers have been extensively studied in the literature [13]. In this work we focus on the classical graph formulation of Ramsey numbers, where $\mathbf{R}(r,s)$ is the smallest number $n$ such that every undirected graph $G$ on $n$ vertices has either has a clique of size $r$, or an independent set of size $s$. While tight bounds on $\mathbf{R}(r,s)$ are known for small values of $r$ and $s$ (namely, for $(3,s)$ with $3\leq s\leq 9$ and $(4,s)$ with $s\in\{4,5\}$ [13]), there is a large gap for many other values of $r$ and $s$. In this work, we focus on improving lower bounds on $\mathbf{R}(r,s)$. If we can exhibit a graph with $n$ vertices that has no cliques of size $r$, nor any independent set of size $s$, we establish that $\mathbf{R}(r,s)\geq n+1$.
Except for very small $r,s$ where analytical methods have worked, prior work has used a variety of search algorithms to computationally derive lower bounds on $\mathbf{R}(r,s)$ (by generating graphs without the appropriate cliques and independent sets); some of these algorithms yielded the best known results for many pairs $r,s$ [13]. We invoke AlphaEvolve [14, 12], a code-mutation agent that uses LLMs (Large-Language-Models) to iteratively evolve code-snippets to generate better lower bounds. For each $r,s$, we use AlphaEvolve to generate search algorithms that produce larger and larger graphs without violating the clique and independence set constraints. (The exact experimental setup is a bit more elaborate.) Thus we use AlphaEvolve as a single meta-algorithm to generate search procedures for meeting or beating Ramsey lower bounds for many values of $r$ and $s$. Our work is similar in spirit to the discovery of extremal combinatorial objects in [11, 8].
1.1 Summary of Our Results
In this paper, we focus on small Ramsey numbers [13] where $r,s\leq 22$. Our main results are that we improve $\mathbf{R}(3,13)$ from $60$ [10] to $61$, $\mathbf{R}(3,18)$ from $99$ [4] to $100$, $\mathbf{R}(4,13)$ from $138$ [3] to $139$, $\mathbf{R}(4,14)$ from $147$ [3] to $148$, and $\mathbf{R}(4,15)$ from $158$ [15] to $159$. Furthermore, we match the SoTA on many other cells in Table 1, including all cells for which the exact value of $\mathbf{R}(r,s)$ is known. The search algorithms used to match or improve these 28 bounds can be categorized into four distinct families of initialization; these are summarized in Table 2 and detailed in the Appendix. Note that for some of these cells, prior work only exhibits the graph implying the bound, without giving the search algorithm that discovered the graph [13].
Note: Our constructions for the the cells $\mathbf{R}(3,13),\mathbf{R}(3,18),\mathbf{R}(4,13),\mathbf{R}(4,14)$ and $\mathbf{R}(4,15)$ are made public on Github at this link.
[…]
Table 1: Ramsey Number $R(r,s)$ lower bounds achieved by AlphaEvolve. Highlights indicate performance status: light blue for optimal lower bounds (matching upper bounds); yellow for matching current best-known lower bounds (where the upper bound remains strictly higher); and green for new results that surpass the prior best lower bound. Algorithms are hyperlinked.
2 AlphaEvolve Meta-search, and Discovered Search Algorithms
AlphaEvolve [12, 14] maintains a family of search algorithms (that search for graphs that establish Ramsey lower bounds), using an LLM to evolve them to find better search algorithms. Typically this process uses a synthetic objective function to guide AlphaEvolve, as a proxy for the true objective (the size of the graph being discovered). This detail will become evident in our presentation of Algorithm 1 below.
AlphaEvolve experimental setup: Algorithm 1 details the way this evolution develops search algorithms that try to improve the current SoTA for $\mathbf{R}(r,s)$ (denoted as $n_{\texttt{SoTA}}$ in Algorithm 1). Here are the core ideas:
  • Evolution: Maintain a population $\mathcal{P}$ of search algorithms. Use a performance-based Select function to choose a parent algorithm from $\mathcal{P}$, then prompt an LLM to generate a mutated version $\texttt{p}_{\texttt{new}}$.
  • Execution: Execute $\texttt{p}_{\texttt{new}}$ to get two graphs: a primary graph $G_{1}$ and a larger ”prospect” graph $G_{2}$.
  • Scoring: If $G_{1}$ is valid, it receives a score based on its size, heavily boosted if it is beyond $n_{\texttt{SoTA}}$. $G_{2}$ contributes an additive bonus inversely proportional to its violation count, guiding the search to push boundary limits. Formally, $\mathbb{E}_{viol}$ represents the expected total number of $r$-cliques and $s$-independent sets in a random graph of size $v_{2}$ (with edge probability 0.5). The score $S$ is updated with a term that increases linearly as the number of violations in $G_{2}$ decreases relative to the expected baseline $\mathbb{E}_{viol}$.
A central feature of Algorithm 1 is that it initializes with an empty baseline graph ($\texttt{p}_{\texttt{base}}$). AlphaEvolve then generates search programs $\texttt{p}_{\texttt{new}}$ aimed at discovering larger graphs. This approach encourages broad exploration, allowing the evolutionary process to learn how to modify and extend smaller, valid lower-bound constructions into larger ones. This is in contrast to an approach where we are specifically searching for graph constructions of a particular size [4, 3, 6].
Input: Ramsey parameters $(r,s,n_{\texttt{SoTA}})$, where $r$ represents the clique size, and $s$ represents the independent set size, and $n_{\texttt{SoTA}}$ is the current best graph size for $\mathbf{R}(r,s)$ from [13]
Output: Program $p^{*}$ with best graph
1
2Init: Population $\mathcal{P}\leftarrow\{\texttt{p}_{\texttt{base}}\}$ // $\texttt{p}_{\texttt{base}}$ returns empty graph
3while compute budget remaining do
4 $\texttt{p}_{\texttt{new}}\leftarrow\texttt{LLM\_Mutation}(\texttt{Select}(\mathcal{P}))$
5 $(G_{1},G_{2})\leftarrow\texttt{p}_{\texttt{new}}.\texttt{run}()$
6 if $G_{1}$ has no $r$-cliques or $s$-independent sets then
7    $v_{1},v_{2}\leftarrow|V(G_{1})|,|V(G_{2})|$
8    $S\leftarrow\begin{cases}4\cdot v_{1}&\text{if }v_{1}>n_{\texttt{SoTA}}\\ 2\cdot v_{1}&\text{if }v_{1}=n_{\texttt{SoTA}}\\ v_{1}&\text{otherwise}\end{cases}$
    // Bonus: Reward low-violation "prospect" graphs
9    if $v_{2}>v_{1}$ then
10       $\mathbb{E}_{viol}\leftarrow\frac{\binom{v_{2}}{r}}{2^{\binom{r}{2}}}+\frac{\binom{v_{2}}{s}}{2^{\binom{s}{2}}}$
11       $S\leftarrow S+\frac{1}{2}\max\left(0,1-\frac{\texttt{count\_viol}(G_{2})}{\mathbb{E}_{viol}}\right)$  where count_viol returns the number of $r$-cliques and $s$-independent sets in the graph
12
13 else
14    $S\leftarrow-1$
15 $\texttt{score}(\texttt{p}_{\texttt{new}})\leftarrow S$
16 $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\texttt{p}_{\texttt{new}},S)\}$
return $\arg\max\limits_{\texttt{p}\in\mathcal{P}}\texttt{score}(\texttt{p})$
Algorithm 1 AlphaEvolve: Ramsey Program Search
Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper and Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper detail the five search algorithms that AlphaEvolve found for the cells $\mathbf{R}(3,13)$, $\mathbf{R}(3,18)$, $\mathbf{R}(4,13)$, $\mathbf{R}(4,14)$, and $\mathbf{R}(4,15)$ respectively. The detailed algorithms for all other colored cells in Table 1 are presented in Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper. Note that each of the search algorithms seems specific to the corresponding cells; their success does not seem to transfer between cells, even when run for a reasonably long time. In Table 2, we summarize the initialization strategies used in the search algorithms. We left out algorithms for $\mathbf{R}(3,3)$ and $\mathbf{R}(3,4)$ because AlphaEvolve just quoted the corresponding adjacency matrices from the LLM’s memory. The search algorithms for each of the cells (while having some differences) are different variants of stochastic search.
If we take a closer look at the search algorithms discovered by AlphaEvolve for the colored cells in Table 1, a pattern starts emerging: a) For some of the cells, a standard (stochastic) search from a generic starting point of Erdős-Rényi graphs suffices, b) For some, it was necessary to seed with specific algebraic graphs (e.g., Paley/cyclic), and c) For some the search had to be guided within some structural restrictions (e.g., staying within the family of cyclic graphs). While clique and independent set counting is often a computational bottleneck in Ramsey constructions, the graph sizes for the instances in Table 1 were relatively small. Thus, clique counting was not a limiting factor for these results.
Comment on the use of AI summarization of the AlphaEvolve generated code: We utilized Gemini 3 to summarize the AlphaEvolve-generated code for all cells and to construct the classifications in Table 2. We scrutinized the underlying code to ensure that these AI-generated summaries reflect the search algorithms. In Table 2 we see a bucketing of the various algorithms by Gemini, with the split being on the method by which each algorithm initialized its graph search.
Random & Stochastic Initialization Paley & Algebraic Seeding Circulant & Cyclic Bootstrap Hybrid, Fractal & Spectral Seeding
R(3,5), R(3,6), R(3,7), R(3,8), R(4,5) R(3,6), R(3,10), R(3,20),
R(4,4), R(4,12), R(4,13),
R(4,15), R(6,7)
R(3,9), R(3,13), R(3,16),
R(3,17), R(3,18), R(3,19),
R(3,21), R(3,22), R(4,6),
R(4,7), R(4,10), R(4,14),
R(5,7), R(5,8), R(5,9)
R(3,6), R(3,11)
Initialization: Random graphs ($G(n,p)$) or empty/greedy baseline. Initialization: Explicit seeding with Paley graphs, cubic, and quadratic residue graphs. Initialization: Bootstrapped from circulant graphs, sum-free sets, or cyclic constructions. Initialization: Complex mixtures: Fractal/self-similar construction, spectral properties, or hybrid pools.
Table 2: Taxonomy of Ramsey Search Algorithms by Initialization Method
3 Comparison to Prior Work
We compare our results against prior work that established the previous state-of-the-art (SoTA) for the entries in Table 1. Our primary reference for these baselines is [13]. Since many historical results rely on personal communications without published algorithmic details, we restrict our comparison to cases where: a) a public reference is available, and b) a computational approach was employed to achieve the lower bound.
  • $\mathbf{R}(4,13),\mathbf{R}(4,14),\mathbf{R}(4,15)$, and $\mathbf{R}(6,7)$: Except for $\mathbf{R}(4,15)$, the previous SoTA for these cells were established in [3]. For $\mathbf{R}(4,15)$, the prior SoTA was established in [15]. We improved the bound for $\mathbf{R}(4,13)$ (and $\mathbf{R}(4,14)$), while matching the bound for $\mathbf{R}(6,7)$. Notably, the search algorithms discovered by AlphaEvolve for $\mathbf{R}(4,13)$ and $\mathbf{R}(6,7)$ (Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper and Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) initialize with the same algebraic structures—cubic residue and Paley graphs, respectively—as those used in [3]. For $\mathbf{R}(4,14)$, Exoo [3] initialized the search with a cubic graph, whereas AlphaEvolve (Algorithm Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) used a class of circulant graphs with varying periodicities. The subsequent optimization strategies for all the cells differ significantly: AlphaEvolve utilizes evolved heuristics rather than the standard simulated annealing employed in [3].
    Our approach for $\mathbf{R}(4,15)$ (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) diverges from the Cayley graph initialization of [3] (which is the SoTA prior to [15]), starting instead from a Paley graph. We unfortunately could not compare to [15] as it is referred to as personal communication in [13].
  • $\mathbf{R}(4,6)$: We match the SoTA for this cell, which was achieved by [5]. While [5] initializes the search with a random graph, AlphaEvolve opts for a cyclic graph initialization. Despite this difference, the subsequent search procedure discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) bears structural similarities to the heuristic search described in [5].
  • $\mathbf{R}(4,10)$: The current SoTA for this cell is attributed to [9], who employed a branch-and-bound search on circulant colorings. In contrast, the algorithm discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) initiates by sampling from a distribution over circulant graphs, followed by a custom local search that integrates sophisticated tabu search mechanisms with sequential growth.
This comparison demonstrates that for many cells — even those where AlphaEvolve only matched the SoTA — AlphaEvolve’s search algorithms differ materially from prior methods, despite often sharing initialization families (e.g., algebraic or cyclic graphs). At a meta-level, the algorithms generated by AlphaEvolve exhibit three salient properties: a) they exploit known algebraic structures relevant to the specific cell; b) unlike many human-designed algorithms, they typically chain multiple heuristics or execute them in parallel; and c) they incorporate custom heuristics for approximate clique and independent set counting to accelerate candidate evaluation. Furthermore, for cases such as $\mathbf{R}(4,10)$ (see Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper), AlphaEvolve’s search strategies appear novel and, to the best of our knowledge, are absent from existing literature.
Prior work on AI and extremal combinatorics: Recent work has studied the use of AI in extremal combinatorics: finding counterexamples to graph theory conjectures [16], the capset problem [14], constructing maximum grid subsets avoiding isosceles triangles [2], discovering extremal Ramanujan graphs [11], and improving bounds for the finite field Kakeya problem [8]. Apart from [16], all the the other results (including ours) have used some variant of AlphaEvolve to discover algorithms that eventually search for extremal combinatorial structures. This contrasts with some other advancements in the space of AI and math/theoretical computer science [1, 17] where new theorems were discovered by directly prompting LLMs (large-language-models). It is unclear at this point whether directly prompting an LLM will result in discovering extremal combinatorial structures.
4 Conclusion
LLM-based systems like AlphaEvolve have been applied to a variety of problems in mathematics and computer science (see [11] for a detailed survey). The area of using AI for Math/CS research is a fast evolving landscape. We approach it from the point of view of making progress on classic and well-studied problems, to derive results that stand the test of time.
Our work lies within the domain of extremal combinatorics and is most closely related to [11], which generated gadgets and extremal graphs to establish new hardness of approximation results. However, while [11] required significant human effort to design proof structures amenable to AlphaEvolve, here the human effort is in crafting the overarching Algorithm 1. Unlike prior computational attempts at improving Ramsey lower bounds [13] — which typically rely on executing bespoke, human-designed heuristics — AlphaEvolve automates the heuristic discovery process: it searches for novel search algorithms.
In this work, we focus on Ramsey lower bounds, where the objective is to generate a witness graph of a target size that avoids specific cliques and independent sets. The complementary challenge of improving Ramsey upper bounds requires a fundamentally different approach: demonstrating the non-existence of any such graph larger than a specific threshold. Unlike lower bound searches, this cannot be solved by constructing a single example. Recently, formal methods have been successfully employed to prove that $\mathbf{R}(4,5)=25$ [7].